OpenStudy (anonymous):

For the function f(x)=1/3x−4, find the value of x for which f(x)=x

OpenStudy (anonymous):

plz help asap!!!

OpenStudy (zzr0ck3r):

y = (1/3)x-4 so x = 3(y+4) f^(-1)(x) = (3(x+4)) f(f^(-1)(x)) = f(3(x+4)) = (1/3)(3(x+4)) -4 = (1/3)(3x+12)-4 = x+4-4 = x

OpenStudy (zzr0ck3r):

or (1/3)x - 4 = x (1/3)x - x = 4 x((1/3)-1) = 4 x = 4/(-2/3) = -6

OpenStudy (anonymous):

-6 is the answer thank you

OpenStudy (zzr0ck3r):

yeah I read it wrong the first time.